The length of a rectangle is 3 in. more than its width. Ther perimeter of the rectangle is 30 in.
a) Define a variable for the width:
b) Write an expression for the length, in terms of width:
c) Write an equation to find the width of the rectangle:
d) What is the width of the rectangle?
e) What is the length of the rectangle?

Answer by Boreal(15235)

width is w
length is w+3
l+w=15, the half perimeter (2l+2w=30)
w+w=3=15
2w+3=15
2w=12
w=6 inches width
w+3=9 inches length